40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jun 25 15:32:16 2019, gautam, Update, IOO, PMC and IMC locked again, some MEDM maintenance PMCctrl.png
    Reply  Wed Jun 26 10:55:13 2019, aaron, Update, IOO, PMC and IMC locked again, some MEDM maintenance 
       Reply  Wed Jun 26 11:11:40 2019, Milind, Update, IOO, PMC and IMC locked again, some MEDM maintenance burtgooey.pdf
    Reply  Wed Jun 26 18:28:24 2019, rana, Update, IOO, PMC and IMC locked again, some MEDM maintenance 
       Reply  Sun Jun 30 22:21:07 2019, Milind, Update, IOO, PMC and IMC locked again, some MEDM maintenance 
          Reply  Sun Jun 30 23:52:09 2019, Koji, Update, IOO, PMC and IMC locked again, some MEDM maintenance 
             Reply  Mon Jul 1 11:02:05 2019, Milind, Update, IOO, PMC and IMC locked again, some MEDM maintenance 
          Reply  Tue Jul 9 10:37:42 2019, Milind, Update, IOO, keyed psl crate, unstick.py, pmc autolocker code- working  
             Reply  Mon Jul 15 14:53:40 2019, Milind, Update, IOO, keyed psl crate, unstick.py 
                Reply  Mon Jul 15 18:55:05 2019, gautam, Update, IOO, Megatron hard-rebooted 
             Reply  Wed Jul 24 12:24:43 2019, Milind, Update, IOO, unstick.py and ifotest 
Message ID: 14696     Entry time: Tue Jun 25 15:32:16 2019     Reply to this: 14699   14701
Author: gautam 
Type: Update 
Category: IOO 
Subject: PMC and IMC locked again, some MEDM maintenance 

Aaron complained to me earlier that the PMC could not be locked. Turned out to be a classic sticky slider problem, I keyed the c1psl VME crate, and did the usual burtrestore trick. After that, I could immediately lock the PMC and IMC with the nominal gain settings.

I also looked at the wiring at the rack. An SLP250 was installed at the mixer IF output, in parallel with a 50 ohm terminator to ground. I removed this, because as Aaron pointed out, the PMC servo card "FP1 TEST" input is already 50 ohm, and has two cascaded LC filter stages immediately after to filter out the 2f component, so the extra low-pass filtering is superfluous (in any case, 250 MHz is much too high a cutoff to be using for cutting out the 2f component which will be at ~70 MHz).

Finally, in the last ~2 weeks, we have been running with the PMC servo gain of +17 (as opposed to +18 from before). The old gain is too high, as noted by Milind. But the MEDM field for this gain goes RED unless the gain is +18. I adjusted the value of the  C1:PSL-STAT_PMC_NOM_GAIN channel to +17, so that this is no longer the case. I also edited the PMC MEDM screen to get rid of my comment that the "SLOW ADC IS DEAD" for the PMC TRANS field, since I have now hooked up the PMC trans photodiode to our temporary Acromag box.

Attachment 1: PMCctrl.png  13 kB  Uploaded Tue Jun 25 16:50:11 2019  | Hide | Hide all
PMCctrl.png
ELOG V3.1.3-